# The synthetic instrument: From sparse association to sparse causation

### v2

@@ 4.3 Extension to nonlinear settings
assuming $\beta$ is set to its true value. Following this, we define the population GMM loss:
@@
For the function $f(X;\beta)$, let $\partial f(X;\beta)/\partial\beta$ denote the $p\times 1$ column vector $\left({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}}\right)^{\mathrm{\scriptscriptstyle T}}$ and let $\partial f(X;\beta)/\partial\beta^{\mathrm{\scriptscriptstyle T}}$ denote the $1\times p$ row vector $\left({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}}\right)$. We make the following assumptions, denoted as D1 and D2, which are generalizations of conditions $A1$ and $A4$ in the nonlinear setting.
@@
    (Nonlinear plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$ and $\dot{\beta}_{C^{*}}\neq 0$, and let the synthetic GMM obtained by assuming $\beta_{{C^{*}}}=0$ be $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}G(\beta).$ The plurality rule assumes that $\widetilde{\beta}^{{C^{*}}}$ is uniquely defined and that $\max\limits_{\beta\in\mathbb{R}^{p}}|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q.$
@@
Suppose that models (1), (3) and conditions A2, D1, and D2 hold.
@@
    If A3 holds, then $\dot{\beta}$ is identifiable via $\dot{\beta}=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, subject to $||\beta||_{0}<p-q$.
@@
    If A3 fails, then $\dot{\beta}$ is not identifiable, and for all $\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, we have $||\widetilde{\beta}||_{0}\geq p-q$.
We discuss how to estimate $\beta$ in the finite sample setting. We replace the expectation in equation (8) with the empirical average. Specifically, we consider the following loss function:
$G_{n}(\beta)=\left(\frac{1}{n}\sum^{n}_{i=1}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right)^{\mathrm{\scriptscriptstyle T}}W\left(\frac{1}{n}\sum^{n}_{i=1}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right),$
where $W\in\mathbb{R}^{(p-q)\times(p-q)}$ is a weight matrix. The oracle weight matrix that achieves the highest efficiency in GMM is the inverse of $\mathbb{E}\left[\text{SIV}(Y-f(X;\dot{\beta}))^{2}\text{SIV}^{\mathrm{\scriptscriptstyle T}}\right]=\widetilde{\sigma}^{2}\text{Cov}(\text{SIV})$, where $\widetilde{\sigma}^{2}=\text{Var}(g(U)+\epsilon_{y})$ (Burgess et al., 2017). In practice, we estimate the weight matrix $W$ using the inverse of empirical covariance: $\widehat{\text{Cov}}^{-1}(\widehat{SIV})$.
Finally, let $\bm{X}\in\mathbb{R}^{n\times p}$, $\bm{Y}\in\mathbb{R}^{n\times 1}$, $f(\bm{X};\beta)\in\mathbb{R}^{n\times 1}$, and $\bm{SIV}\in\mathbb{R}^{n\times(p-q)}$ denote the relevant random matrices in the finite sample setting. We consider the following optimization problem:
@@
The unconstrained optimization in (9) is also referred to as the nonlinear two-stage least squares estimator in the literature (Amemiya, 1974). Under the linear setting where $f(\bm{X};\beta)=\bm{X}\beta$, (9) has a similar form to (7). Optimization (9) can be accelerated using a splicing approach, which is designed to efficiently solve the best subset selection problem (Zhu et al., 2020; Zhang et al., 2023).
@@ 5 Simulation studies
In this section, we evaluate the numerical performance of the proposed SIV method and compare it with other methods across various scenarios. First, we assess the performance of the SIV method under a linear outcome model in Section 5.1. Next, we evaluate the algorithm in the context of a nonlinear outcome model in Section 5.2. In Section LABEL:sec:MS of the supplementary material, we explore an alternative moment selection estimator inspired by the work of Andrews (1999) and compare it with our proposed estimator.
@@ 5.1 Simulation studies with a linear outcome model
We begin by evaluating the SIV estimator within a linear outcome model. The model is defined by $f(X;\beta)=X^{\mathrm{\scriptscriptstyle T}}\beta$ and $g(U)=U^{\mathrm{\scriptscriptstyle T}}\gamma$. In our simulations, we let $q=3$, $s=5$, and $\beta=(1,1,1,1,1,0,\ldots,0)^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{p}$. Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\text{Uniform}(-1,1)$ for $j=1,\ldots,p$, and $k=1,\ldots,q$. The hidden variables $U_{i,k}$ follow i.i.d. standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$. The random errors are generated as $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=5$. We evaluate the performance of our method under the following two settings: (i) Low-dimensional cases: $p=100$ and $n\in\{200,600,1000,\ldots,5000\}$; (ii) High-dimensional cases: $n=500$ and $p\in\{500,750,1000,\ldots,3000\}$. All simulation results are based on 1000 Monte Carlo runs. The data-generating mechanism is designed to mimic certain features of the real application in Section 6; see Section LABEL:sec:comparison_sim&data in the supplementary material for more details.
@@
    (Trim, Ćevid et al., 2020): We implement Ćevid et al. (2020)’s method using the code available from https://github.com/zijguo/Doubly-Debiased-Lasso.
@@
    (Null, Miao et al., 2023): For the low-dimensional settings, we implement Miao et al. (2023)’s method using the code available from https://www.tandfonline.com/doi/suppl/10.1080/01621459.2021.2023551. For the high-dimensional settings, their method cannot be applied directly because $\xi$ in their method cannot be solved by ordinary least squares. Therefore, we replace ordinary least squares with the lasso, where the tuning parameter is selected by 10-fold cross-validation.
@@
    (IV-Lasso): Motivated by a reviewer’s comment, we consider the following “IV-Lasso” estimator. In the first step, we use Lasso to select active predictors and define $\widehat{\mathcal{A}}=\{j:\widetilde{\beta}_{j}\neq 0\}$, where $\widetilde{\beta}$ is the Lasso estimator described above. In the next step, we use $\widehat{SIV}$ as instrumental variables to estimate the treatment effects of $X_{\widehat{\mathcal{A}}}$ on $Y$ using the ivreg package, denoted as $\widehat{\beta}^{\;\text{IV-Lasso}}_{\widehat{\mathcal{A}}}$, and set $\widehat{\beta}^{\;\text{IV-Lasso}}_{\widehat{\mathcal{A}}^{c}}=0$.
@@
We present the $\ell_{1}$-estimation errors of the five methods in Figures 2. In low-dimensional cases, as illustrated in Figure 2(a), the bias of the Lasso estimator stabilizes as the sample size grows. The Lasso estimator does not account for unmeasured confounding and is, therefore, not expected to perform well. The Lasso, Trim, and Null methods exhibit considerable $\ell_{1}$-estimation bias relative to other methods. Compared to IV-Lasso, our SIV estimator demonstrates superior performance. This advantage arises from the $\ell_{0}$-optimization employed in the SIV method, which enables more accurate variable selection, as illustrated in Figure 3(a) later.
@@
For the high-dimensional settings, we can see from Figure 2(b) that our SIV method consistently outperforms the comparison methods. As discussed in Section 1, the true correlations between $X$ and $Y$ are non-sparse. This explains the large estimation errors of the Lasso method. The Null method exhibits even larger biases than the Lasso method. The Trim method, designed specifically for high-dimensional settings, outperforms both the Lasso and Null methods. However, our estimator still shows a much smaller bias compared to the Trim estimator.
@@
We also observe from Figure 2(b) that the IV-Lasso estimator underperforms compared to the naive Lasso estimator in high-dimensional settings. This underperformance occurs because the naive Lasso method applies $\ell_{1}$ penalization, which drives the coefficients of incorrectly selected $\widehat{\beta}_{i}$ towards zero. In contrast, the second step of IV-Lasso negates this shrinkage, causing the coefficients of incorrectly selected $\widehat{\beta}_{i}$ to diverge from zero, thereby increasing estimation bias.
[…]
@@
We further evaluate the performance of our proposed algorithm in settings with non-diagonal $\text{Cov}(\epsilon_{x})$. We generate $\epsilon_{x,i}\sim\mathbb{N}(0,D)$. For the low-dimensional setting, we randomly select 20 pairs of $i,j\in\{1,2,\ldots,p\}$ and set $D_{i,j}=D_{j,i}=1$. The list of pairs is provided in Section LABEL:sec:detail:nondiagonal of supplementary material. In the high-dimensional settings, we set $D_{i,j}=4\times 0.3^{|i-j|}$. All other aspects of the data-generating mechanism remain unchanged. In the low-dimensional scenario, we use the Robust Principal Component Analysis method (Candès et al., 2011) to estimate the low-dimensional structure of the covariance matrix and factor loadings. For the high-dimensional scenario, we directly use the Principal Component Analysis method to estimate factor loadings. These are implemented using the R packages rpca and prcomp, respectively.
The simulation results, presented in Figures 4 and 5
@@
illustrate the $\ell_{1}$-estimation errors and false discovery rates for various methods. These methods exhibit similar trends in $\ell_{1}$-error performance as observed in the previous setting with uncorrelated errors. The Trim method shows a slight improvement, achieving lower $\ell_{1}$-errors and false discovery rates compared to the previous simulation results. The performance of the SIV method remains consistent with its performance when $D$ is a diagonal matrix. Additionally, our method continues to outperform the other comparison methods.
@@ 5.2 Simulation studies with non-linear outcome models
We then evaluate the performance of our proposed estimator (9) with nonlinear outcome models. We set $q=2$, $s=2$, and $p=10$. We consider two different settings for $f(X)$. In the first setting, $f(X;\beta)=\sum^{10}_{j=1}X_{j}^{3}\beta_{j}$ with $\beta=(0.3,0.3,0,0,\ldots,0)^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{10}$ and $g(U)=U_{1}^{3}\gamma_{1}+U_{2}^{3}\gamma_{2}$. In the second setting, $f(X;\beta)=\exp(X^{\mathrm{\scriptscriptstyle T}}\beta)$ and $g(U)=(U^{3})^{{\mathrm{\scriptscriptstyle T}}}\gamma$. Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\mathbb{N}(0,1)$ for $j=1,\ldots,p$ and $k=1,\ldots,q$. The hidden variables $U_{i,k}$ follow i.i.d. standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$. The random errors are generated by $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=1$. We evaluate the performance of our estimators with $n\in\{1000,\ldots,5000\}$. All simulation results are based on 1000 Monte Carlo runs.
For comparison, we did not implement the other methods in Section 5.1, as they are not designed for non-linear outcome models. Instead, we consider a popular method for addressing endogeneity in high-dimensional settings (e.g. Wang and Blei, 2019; Ouyang et al., 2023; Fan et al., 2024), which first estimates the unmeasured confounders $U$ and then directly adjusts for its estimate $\widehat{U}$ in the regression modeling. Specifically, we compare the proposed method with the following two variations of the so-called U-hat method, with the tuning parameter $k$ selected using 10-fold cross-validation in all cases:
@@
    (U-hat1): First, we obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ using the equation $\widehat{\bm{U}}=\bm{X}\widehat{\text{Cov}}(X)^{-1}\widehat{\Lambda}$. Next, we obtain $\widehat{\beta}$ by solving the following optimization problem: $\widehat{\beta}=\underset{\beta\in\mathbb{R}^{p},\gamma\in\mathbb{R}^{q}}{\operatorname*{arg\,min}}\|\bm{Y}-f(\bm{X};\beta)-\widehat{\bm{U}}\gamma\|_{2}^{2}\quad\text{subject to}\quad\|\beta\|_{0}\leq k.$
@@
    (U-hat2): We first obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ similarly. We then obtain $\widehat{\beta}$ by solving the following optimization problem: $\widehat{\beta}=\underset{\beta\in\mathbb{R}^{p},\gamma\in\mathbb{R}^{q}}{\operatorname*{arg\,min}}\|\bm{Y}-f(\bm{X};\beta)-\widehat{\bm{U}}^{3}\gamma\|_{2}^{2}\quad\text{subject to}\quad\|\beta\|_{0}\leq k,$ where $\widehat{\bm{U}}^{3}\in\mathbb{R}^{n\times q}$ is defined as $\{\widehat{\bm{U}}^{3}\}_{i,j}=\{\widehat{\bm{U}}_{i,j}\}^{3}$. Note this method assumes knowledge of the specific form of $g(U)$, which is generally not available in practice.
Figure 6 graphically presents the $\ell_{1}$-estimation error for all estimators. The results show that U-hat1 and U-hat2 have similar performance, with bias stabilizing as the sample size increases. In contrast, the bias of our estimator decreases with larger samples, indicating that our method consistently estimates the causal parameter $\beta$ in nonlinear outcome models.
@@
Under the linear models in Section 5.1, results from the U-hat (including U-hat1 and U-hat2) and SIV methods coincide numerically. While the $U$-hat method has been proposed in the literature (e.g., Wang and Blei, 2019), its validity has been widely challenged (e.g., Ogburn et al., 2019; Grimmer et al., 2023). From this perspective, our results may be viewed as a justification for the $U$-hat method in the special and unrealistic case of linear outcome model (2). In general, however, our method is fundamentally different from the U-hat method.
The U-hat method relies on regression adjustment, which assumes a specified outcome-confounder relationship. If confounders $U$ were observed, then, at least in principle, this relationship could be estimated from the data. But if $U$ is unobserved, then specifying the functional form of $g$ depends entirely on subject-matter knowledge, which is rarely available. In Section LABEL:sec:Uhat of the supplementary material, we explain in detail why the $U$-hat method fails under the general model $Y=f(X;\beta)+g(U)+\epsilon_{y}$.
In contrast, our approach constructs instrumental variables, which are agnostic to the form of unmeasured confounding even if the causal relationship is non-linear. Because of this, we believe it is a more robust approach in this setting. To our knowledge, we are the first to apply the IV framework to this problem, offering a method that accommodates nonlinear causal effects and arbitrary dependence on unmeasured confounders $U$. See Section LABEL:sec:Uhat in the supplementary materials for more discussions.

### v3

@@ 4.3 Extension to nonlinear settings
when $\beta$ is set to its true value. Following this, we define the population GMM loss:
@@
For the function $f(X;\beta)$, let $\partial f(X;\beta)/\partial\beta$ denote the $p\times 1$ column vector $({\partial f(X;\beta)}/{\partial\beta_{1}}$ $\ldots,{\partial f(X;\beta)}/{\partial\beta_{p}})^{\mathrm{\scriptscriptstyle T}},$ and let $\partial f(X;\beta)/\partial\beta^{\mathrm{\scriptscriptstyle T}}$ denote the $1\times p$ row vector $({\partial f(X;\beta)}/{\partial\beta_{1}},\ldots,$ ${\partial f(X;\beta)}/{\partial\beta_{p}})$. We make the following assumptions, denoted as D1 and D2, which are generalizations of Conditions A1 and A4 in the nonlinear setting.
@@
    (Nonlinear plurality rule) Let $C^{*}$ be a subset of $\{1,2,\ldots,p\}$ with cardinality $q$ and $\dot{\beta}_{C^{*}}\neq 0$, and let the synthetic GMM estimator obtained by assuming $\beta_{C^{*}}=0$ be $\widetilde{\beta}^{C^{*}}=\underset{\beta\in\mathbb{R}^{p}:\beta_{C^{*}}=0}{\operatorname*{arg\,min}}G(\beta).$ The plurality rule assumes that $\widetilde{\beta}^{C^{*}}$ is uniquely defined and that $\max\limits_{\beta\in\mathbb{R}^{p}}|\{C^{*}:\widetilde{\beta}^{C^{*}}=\beta\}|\leq q.$
@@
Suppose that models (1), (3), and Conditions A2, D1, and D2 hold.
@@
    If A3 holds, then $\dot{\beta}$ is identifiable via $\dot{\beta}=\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, subject to $\|\beta\|_{0}<p-q$.
@@
    If A3 fails, then $\dot{\beta}$ is not identifiable, and for all $\widetilde{\beta}\in\underset{{\beta}\in\mathbb{R}^{p}}{\operatorname*{arg\,min}}\;G(\beta)$, we have $\|\widetilde{\beta}\|_{0}\geq p-q$.
We now discuss the estimation of $\beta$ in finite samples. A natural approach is to replace the expectation in equation (9) with its empirical counterpart. Specifically, we consider the following loss function:
$G_{n}(\beta)=\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right)^{\mathrm{\scriptscriptstyle T}}W\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right),$
where $W\in\mathbb{R}^{(p-q)\times(p-q)}$ is a weight matrix.
We now discuss the estimation of $\beta$ in finite samples. A natural approach is to replace the expectation in equation (9) with its empirical counterpart. Specifically, we consider the following loss function:
$G_{n}(\beta)=\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right)^{\mathrm{\scriptscriptstyle T}}W\left(\frac{1}{n}\sum_{i=1}^{n}[SIV_{i}\{Y_{i}-f(X_{i};\beta)\}]\right),$
[…]
Finally, let $\bm{X}\in\mathbb{R}^{n\times p}$, $\bm{Y}\in\mathbb{R}^{n\times 1}$, $f(\bm{X};\beta)\in\mathbb{R}^{n\times 1}$, and $\bm{SIV}\in\mathbb{R}^{n\times(p-q)}$ denote the relevant random matrices in the finite-sample setting. In the low-dimensional setting, where the number of instruments pqp-q is fixed and smaller than $n$, we consider the following optimization problem:
@@
The unconstrained optimization in (10) is also referred to as the nonlinear two-stage least squares estimator in the literature (Amemiya, 1974). Under the linear setting where $f(\bm{X};\beta)=\bm{X}\beta$, equation (10) has a similar form to (7). Optimization (10) can be accelerated using a splicing approach, which is designed to efficiently solve the best subset selection problem (Zhu et al., 2020; Zhang et al., 2023). We establish the theoretical properties of (10) in Section LABEL:sec:appthmnonlinear of the supplementary material, showing that under mild regularity conditions, the proposed estimator achieves both consistency and variable selection consistency in the low-dimensional setting.
Remark 4.
In high-dimensional settings where the dimension of $\operatorname{SIV}$, namely pqp-q, exceeds the sample size $n$, the matrix $\operatorname{SIV}^{\mathrm{\scriptscriptstyle T}}\operatorname{SIV}$ is not invertible. In this case, one may follow the approach of Belloni et al. (2018, p. 35) and choose the weight matrix $W$ as
$W=\operatorname{diag}\!\left(\operatorname{Var}\!\left(\mathbb{E}_{n}\!\left[\operatorname{SIV}(Y-f(X;\widetilde{\beta}))\right]\right)\right)^{-1},$
where $\widetilde{\beta}$ denotes an initial estimator of $\beta$.
@@ 5 Simulation studies
In this section, we evaluate the numerical performance of the proposed SIV method and compare it with other methods across various scenarios. First, we assess the performance of the SIV method under a linear outcome model in Section 5.1. Next, we evaluate the algorithm in the context of a nonlinear outcome model in Section 5.2.
We provide additional simulation results in the supplementary material. In Section LABEL:sec:simuweakeffect, we present simulation results for various estimators under dense confounding with many weak effects. In Section LABEL:sec:MS, we explore an alternative moment selection estimator, inspired by the work of Andrews (1999), and compare it with our proposed estimator. Section LABEL:sec:CI_sup discusses the construction of confidence intervals for selected causal variables using the ivreg function. Finally, in Section LABEL:sec:countdata, we extend the SIV algorithm to confounded count data, where the effect of the unmeasured confounder on the outcome cannot be additively separated from the effect of the treatment.
@@ 5.1 Simulation studies with a linear outcome model
We begin by evaluating the SIV estimator within a linear outcome model. The model is defined by $f(X;\beta)=X^{\mathrm{\scriptscriptstyle T}}\beta$ and $g(U)=U^{\mathrm{\scriptscriptstyle T}}\gamma$. In our simulations, we let $q=3$, $s=5$, and $\beta=(1,1,1,1,1,0,\ldots,0)^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{p}$. Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\text{Uniform}(-1,1)$ for $j=1,\ldots,p$ and $k=1,\ldots,q$. The hidden variables $U_{i,k}$ follow i.i.d. standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$. The random errors are generated as $\epsilon_{x}\sim\mathbb{N}(0,\sigma_{x}^{2}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=5$. We evaluate the performance of our method under the following two settings: (i) low-dimensional cases: $p=100$ and $n\in\{200,600,1000,\ldots,5000\}$; (ii) high-dimensional cases: $n=500$ and $p\in\{500,750,1000,\ldots,3000\}$. All simulation results are based on $1000$ Monte Carlo runs. The data-generating mechanism is designed to mimic key features of the real application in Section 6; see Section LABEL:sec:comparison_sim&data of the supplementary material for a comparison between the real and synthetic data.
@@
    (Trim, Ćevid et al., 2020): We implement Ćevid et al. (2020)’s method using the code available from https://github.com/zijguo/Doubly-Debiased-Lasso, with an update detailed in Section LABEL:sec:update_supp of the supplementary material.
@@
    (Null, Miao et al., 2023): For the low-dimensional settings, we implement Miao et al. (2023)’s method using the code available from https://www.tandfonline.com/doi/suppl/10.1080/01621459.2021.2023551. For the high-dimensional settings, their method cannot be applied directly because $\xi$ in their procedure cannot be solved by ordinary least squares. Therefore, we replace ordinary least squares with the Lasso, with the tuning parameter selected by 10-fold cross-validation.
@@
    (IV-Lasso) Motivated by a reviewer’s suggestion, we consider the following two-step “IV-Lasso” procedure. First, we apply the Lasso as a screening step to identify candidate causal predictors of $Y$. Specifically, we solve
    $\widetilde{\beta}=\underset{\beta\in\mathbb{R}^{p}}{\arg\min}\left\{\|{\bf Y}-\widehat{{\bf X}}\beta\|_{2}^{2}+\lambda\|\beta\|_{1}\right\},$
[…]
@@
We present the $\ell_{1}$-estimation errors of the five methods in Figures 2. In low-dimensional cases, as illustrated in Figure 2(a), the bias of the Lasso estimator stabilizes as the sample size grows. The Lasso estimator does not account for unmeasured confounding and is therefore not expected to perform well. The Lasso, Trim, and Null methods exhibit considerable $\ell_{1}$-estimation bias relative to other methods. Compared to IV-Lasso, our SIV estimator demonstrates superior performance. This advantage arises from the $\ell_{0}$-optimization employed in the SIV method, which enables more accurate variable selection, as illustrated in Figure 3(a) later.
To ensure a fair comparison across all methods, we use cross-validation to select $\lambda$. However, it is well known that cross-validation can lead to overfitting and an increased false discovery rate. In practice, researchers may adopt the “one-standard-error” rule, which selects the largest $\lambda$ such that the cross-validation error remains within one standard error of its minimum (Hastie et al., 2009; Kang et al., 2016). In the simulation settings we have considered, we find that the “one-standard-error rule” allows the IV-Lasso method to perform as well as the SIV method. See Section LABEL:sec:ivlasso1se of the supplementary material for details on how the “one-standard-error” rule improves the empirical performance of IV-Lasso.
@@
For the high-dimensional settings, we can see from Figure 2(b) that our SIV method consistently outperforms the comparison methods. As discussed in Section 1, the true correlations between $X$ and $Y$ are non-sparse. This explains the large estimation errors of the Lasso method. The Null method exhibits even larger biases than the Lasso method. The Trim method, designed specifically for high-dimensional settings, outperforms both the Lasso and Null methods. However, our estimator still shows a much smaller bias than the Trim estimator. In Section LABEL:sec:trim_sup of the supplementary material, we provide additional discussion, simulations, and comments on why the Trim estimator performs less favorably compared to our estimator.
@@
We also observe from Figure 2(b) that the IV-Lasso estimator underperforms compared to the naive Lasso estimator in high-dimensional settings. This underperformance occurs because the naive Lasso method applies $\ell_{1}$-penalization, which drives the coefficients of incorrectly selected $\widehat{\beta}_{i}$ towards zero. In contrast, the second step of IV-Lasso negates this shrinkage, causing the coefficients of incorrectly selected $\widehat{\beta}_{i}$ to diverge from zero, thereby increasing estimation bias.
Since the underlying $\beta$ is sparse, we also report the performance of variable selection for all the methods in Figure 3. All these methods correctly classify the true causes of the outcome as active exposures, that is, $\widehat{\mathcal{A}}\supseteq\mathcal{A}$. Thus, we only report the average false discovery rates, denoted as $\{\widehat{\mathcal{A}}\setminus\mathcal{A}\}/\#\widehat{\mathcal{A}}$, across 1000 Monte Carlo runs. It is evident that our proposal achieves the lowest false discovery rate among all the methods in both the low- and high-dimensional settings.
@@
We further evaluate the performance of our proposed algorithm in settings with non-diagonal $\text{Cov}(\epsilon_{x})$. We generate $\epsilon_{x,i}\sim\mathbb{N}(0,D)$. For the low-dimensional setting, we randomly select 20 pairs of $i,j\in\{1,2,\ldots,p\}$ and set $D_{i,j}=D_{j,i}=1$. The list of pairs is provided in Section LABEL:sec:detail:nondiagonal of the supplementary material. In high-dimensional settings, we set $D_{i,j}=4\times 0.3^{|i-j|}$. All other aspects of the data-generating mechanism remain unchanged. In the low-dimensional scenario, we use the Robust Principal Component Analysis method (Candès et al., 2011) to estimate the low-dimensional structure of the covariance matrix and factor loadings. For the high-dimensional scenario, we directly use the Principal Component Analysis method to estimate factor loadings. These are implemented using the R packages rpca and prcomp, respectively.
@@
The simulation results, presented in Figures 4 and 5, illustrate the $\ell_{1}$-estimation errors and false discovery rates for various methods. These methods exhibit similar trends in $\ell_{1}$-error performance as observed in the previous setting with uncorrelated errors. The Trim method shows a slight improvement, achieving lower $\ell_{1}$-errors and false discovery rates compared to the previous simulation results. The performance of the SIV method remains consistent with its performance when $D$ is a diagonal matrix. Additionally, our method continues to outperform the other comparison methods.
@@ 5.2 Simulation studies with non-linear outcome models
We then evaluate the performance of our proposed estimator (10) with nonlinear outcome models. We set $q=2$, $s=2$, and $p=10$. We consider two different settings for $f(X)$. In the first setting, $f(X;\beta)=\sum^{10}_{j=1}X_{j}^{3}\beta_{j}$ with $\beta=(0.3,0.3,0,0,\ldots,0)^{\mathrm{\scriptscriptstyle T}}\in\mathbb{R}^{10}$ and $g(U)=U_{1}^{3}\gamma_{1}+U_{2}^{3}\gamma_{2}$. In the second setting, $f(X;\beta)=\exp(X^{\mathrm{\scriptscriptstyle T}}\beta)$ and $g(U)=(U^{3})^{{\mathrm{\scriptscriptstyle T}}}\gamma$. Each element in $\Lambda_{j,k}$ and $\gamma_{k}$ is independently generated from $\mathbb{N}(0,1)$ for $j=1,\ldots,p$ and $k=1,\ldots,q$. The hidden variables $U_{i,k}$ follow i.i.d. standard normal distributions for $i=1,\ldots,n$ and $k=1,\ldots,q$. The random errors are generated as $\epsilon_{x}\sim\mathbb{N}(0,\sigma^{2}_{x}I_{p})$ and $\epsilon_{y}\sim\mathbb{N}(0,\sigma^{2})$, where $\sigma_{x}=2$ and $\sigma=1$. We evaluate the performance of our estimators with $n\in\{1000,\ldots,5000\}$. All simulation results are based on 1000 Monte Carlo runs.
For comparison, we did not implement the other methods in Section 5.1, as they are not designed for nonlinear outcome models. Instead, we consider a popular method for addressing endogeneity in high-dimensional settings (e.g. Wang and Blei, 2019; Ouyang et al., 2023; Fan et al., 2024), which first estimates the unmeasured confounders $U$ and then directly adjusts for its estimate $\widehat{U}$ in the regression modeling. Specifically, we compare the proposed method with the following two variations of the so-called U-hat method, with the tuning parameter $k$ selected using 10-fold cross-validation in all cases:
@@
    (U-hat1): First, we obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ using the equation $\widehat{\bm{U}}=\bm{X}\widehat{\text{Cov}}(X)^{-1}\widehat{\Lambda}$. Next, we obtain $\widehat{\beta}$ by solving the following optimization problem:
[…]
@@
    (U-hat2): We first obtain $\widehat{\bm{U}}\in\mathbb{R}^{n\times q}$ similarly. We then obtain $\widehat{\beta}$ by solving the following optimization problem:
[…]
    where $\widehat{\bm{U}}^{3}\in\mathbb{R}^{n\times q}$ is defined as $\{\widehat{\bm{U}}^{3}\}_{i,j}=\{\widehat{\bm{U}}_{i,j}\}^{3}$. Note that this method assumes knowledge of the specific form of $g(U)$, which is generally not available in practice.
Figure 6 displays the $\ell_{1}$-estimation errors for all estimators. The results indicate that U-hat1 and U-hat2 perform similarly, with their biases stabilizing as the sample size increases. In contrast, the bias of the proposed estimator decreases with larger sample sizes, demonstrating its consistency in estimating the causal parameter $\beta$ under nonlinear outcome models. In Section LABEL:sec:nonlinear,nondigonal of the supplementary material, we further consider a setting with a nonlinear outcome model and a non-diagonal $\text{Cov}(\epsilon_{x})$. The results suggest that the SIV method remains consistent in this complex setting.
@@
Under the linear models in Section 5.1, results from the U-hat (including U-hat1 and U-hat2) and SIV methods coincide numerically. While the $U$-hat method has been proposed in the literature (e.g., Wang and Blei, 2019), its validity has been widely challenged (e.g., Ogburn et al., 2019; Grimmer et al., 2023). From this perspective, our results may be viewed as a justification for the $U$-hat method in the special and unrealistic case of the linear outcome model (2).
In general, our approach differs fundamentally from the U-hat methods. The U-hat1 method yields consistent estimators of $\beta$ only when the treatment–outcome relationship is linear. As shown in Section LABEL:sec:Uhat of the supplementary material, U-hat1 fails under the more general model
$Y=f(X;\beta)+g(U)+\epsilon_{y},$
where $f(X;\beta)$ may be nonlinear. In particular, we derive a necessary condition (Equation (LABEL:eqn:identifiabilityequation_supp)) that must hold for U-hat1 to consistently estimate $\beta$, but this condition is typically violated in nonlinear outcome models. The U-hat2 method, on the other hand, relies on modeling the latent confounder–outcome relationship, introducing additional assumptions that are implausible in practice and do not resolve the underlying identification challenge.
In contrast, our approach constructs instrumental variables, which are agnostic to the form of unmeasured confounding, even if the causal relationship is nonlinear. Because of this, we believe it is a more robust approach in this setting. To our knowledge, we are the first to apply the IV framework to this problem, offering a method that accommodates nonlinear causal effects and arbitrary dependence on unmeasured confounders $U$. See Section LABEL:sec:Uhat of the supplementary material for further discussion and a comparison between the proposed method and the U-hat methods.